2.         The Ka of nitrous acid HNO2is 4.5 x 10-4.  In a titration, 50.0 mL of 1.00 M HNO2is tit  rated with 0.750 M NaOH.     

            a. Calculate the pH of the solution:

                        (i) Before the titration begins.

(ii) When sufficient NaOH has been added to neutralize half the nitrous acid                   originally present.

                        (iii) At the equivalence point.

(iv)When 0.05 mL NaOH less than that required to reach the equivalence point  has been added.

(v) When 0.05 mL NaOH more than that required to reach the equivalence point has been added.

            (b) Can Bromthymol blue be used as the indictor for this titration? Explain your answer.

            (c) Will methyl red be a satisfactory indicator here?

3. A 0.500L solution contains 0.025 mol Ag+.

(a) Calculate the minimum mass of NaCl that must be added to precipitate AgCl from the solution.

(b) If excess Cl-is added to the solution the AgCl precipitate dessolves due to the formation of  [Ag(Cl)2]-= 1.8 x 10 5.  Calculate the minimum amount Cl-that must be added to dissolve the precipitate.

-we are given the weak acid and the conjugate base in the problem. we need to go from there. First find the pKa for the acid then proceed from t

----------

For number 2

Start by writing the balanced chemical equation.

then proceed as we did in the lecture slides to solve for pH.

For number three

this is a Ksp problem

We did one of these on the lecture slides as well

Again start with a balanced equation then write a Ksp expression for the reaction.

yu can then solve for [Cl-]
Answer #1

solution :- Ka of Nitrous Acid = 4.5810-4 ANO2 + Naon Na NO₂ + H2O moles of Naona males of HNO2 MINI = M2 V2 0.750m xvi = 1.0a= -b I ſ b²-4ac 2х4 k -Ч: 5 xo 9 4 2-625 xot-4xi x+(-4.5 xis) 2. ។ x = -4.5 X10 + 2 - 025 Хо +o-ool 2. ч. 5 Хо Ч+ -Оч?ч 2.there will be only At equivalence point NaNO2 in the solution, maley of Nano2 formed = moles of HNO2 50 mmoli Total volume ofSIN 0.05 ml Naon less that that equivalence point; 66.62 ml of Naon is added; V= 66.67 -0.05 = 50 nmol. initial male of NO 2[ Naon] = 0-0375 mmal. 116.72 me. 0.000321 malah 0.000321 mal/ (ou) (44) 1.0 x10.14 0.000321 (H+) (n+) 3.112 x 10 = 3112.53X